OpenStudy (anonymous):

Calculate the concentration of nitric acid in moles per litre in a sample which has a density, 1.41 g mL–1 and the mass per cent of nitric acid in it being 69%. I'm Stuck :(

OpenStudy (cherylim23):

Okay, you know that mass = density x volume and moles = concentration x volume volume = mass/density and volume = moles/concentration Hence mass/density = moles/concentration --(1) Let the mass of sample be x grams. moles = mass/molar mass = x / (1+14 + 16x3) = x/63 Since the answer requires the form of moles per litre, change the unit of density to 1410g/L Now that you have all these information, sub it into equation (1) 0.69x/1410 = (x/63)/c, where c is concentration x is just a dummy variable here, solve the eqn and you should get c=32M (weird how it's so highly concentrated, but I checked through - don't think I made any mistakes)